# 4Diffusion: Multi-view Video Diffusion Model for 4D Generation

### v1

@@ 1 Introduction
In this paper, we present a novel 4D generation pipeline, namely 4Diffusion, to create high-quality spatial-temporally consistent 4D content from a monocular video. Specifically, we propose a unified diffusion model, 4DM, to capture multi-view spatial-temporal correlations for multi-view video generation. This model provides multi-view spatial-temporal guidance for 4D generation. We construct 4DM based on the powerful pre-trained 3D-aware diffusion model [52], which already achieves high-quality multi-view spatial consistency. We seamlessly integrate a motion module into the 3D-aware diffusion model to extend the temporal modeling capability. Contrary to previous attempts [15, 14] that typically demand extensive large-scale video datasets for tuning the motion module, 4DM achieves reasonable temporal consistency and captures multi-view spatial-temporal correlations after training on only hundreds of multi-view videos. Importantly, we keep the parameters of the 3D-aware diffusion model unchanged to preserve the generalizability and spatial consistency. Subsequently, we propose 4D-aware SDS loss to distill prior knowledge from 4DM to optimize 4D representation parameterized by dynamic NeRF. This approach eliminates discrepancies arising from multiple diffusion models and stabilizes the training process. Moreover, we use 4DM to generate anchor videos conditioned on the input monocular video and devise an anchor loss to enhance the appearance details, facilitating the learning of dynamic NeRF. Finally, we generate 4D content with high-quality spatial-temporal consistency and vibrant motion coherence with the input video as shown in Fig. 1. Qualitative and quantitative experiments demonstrate that our method achieves state-of-the-art performance on multi-view video generation and 4D generation from monocular videos.
@@ 2 Related Work
Similar to us, [19, 42, 63, 62, 57] generate 4D content from a monocular video. Consistent4D [19] introduces an interpolation-driven loss between two adjacent frames to enhance spatial-temporal consistency. However, Consistent4D lacks temporal modeling cross frames. DreamGaussian4D [42], 4DGen [62], and SC4D [57] combine 4D Gaussian Splatting [56, 18] into 4D generation pipeline. Although they notably reduce optimization time, they may result in blurred appearance and inaccurate geometry due to the explicit characteristics of Gaussians. STAG4D [63] proposes a training-free strategy to generate sparse anchor multi-view videos for 4D generation. In contrast, we propose a multi-view video diffusion model to provide multi-view spatial-temporal consistency prior to optimize dynamic NeRF.
@@ 3 Method
Given a monocular video with frames and an optional textual caption, our goal is to generate a high-quality spatial-temporally consistent 4D content, capable of rendering from any novel viewpoint across the temporal dimension. In Sec. 3.1, we talk about 3D-aware diffusion models, employed as the initialization of our unified diffusion model. As shown in Fig. 3, we propose a unified diffusion model 4DM to capture multi-view spatial-temporal correlations for multi-view video generation, which is achieved by inserting a learnable motion module into a frozen 3D-aware diffusion model, as described in Sec. 3.2. Subsequently, we elaborate on distilling prior knowledge from 4DM to optimize 4D representation parameterized by dynamic NeRF and devise an anchor loss to enhance the appearance details, as detailed in Sec. 3.3.
3.1 Preliminary: 3D-aware Diffusion Model
@@
Although we can directly use the original ImageDream to generate a set of individual multi-view images to form multi-view videos, the result lacks temporal consistency as ImageDream has no layer for temporal modeling, as shown in Fig. 5. We thus add a zero-initialized motion module at the end of each block of the UViT network. Specifically, we design our motion modules to consist of several self-attention blocks. Then, each attention block of 4DM includes a spatial module and a motion module . The spatial module comprises a 3D self-attention module and a cross-attention module. We first concatenate the monocular video latent and four multi-view video latents encoded by VAE [21] to obtain a batch of latents , where is the number of channels, and are spatial resolutions, is the number of viewpoints, and is the number of frames. Subsequently, we reshape the temporal axis into the batch dimension and independently process multi-view video latents through the 3D self-attention module,
@@
We utilize Objaverse dataset [11] to train 4DM. Although Objaverse provides nearly 44K animated 3D shapes, rendering multi-view videos and training a diffusion model are time- and computation-consuming using the entire dataset. Moreover, it is worth noting that the Objaverse dataset contains a significant amount of flawed data. Consequently, we manually select a curated subset of 926 high-quality animated 3D shapes from Objaverse dataset [11]. We render multi-view videos from those animated 3D shapes to tune our motion module while holding the parameters of the origin ImageDream frozen. Surprisingly, 4DM successfully learns reasonable temporal dynamics and preserves the characteristics of the origin ImageDream model, including spatial consistency and image understanding ability, even when trained on a small curated dataset. As Fig. 5 illustrates, 4DM generates multi-view spatial-temporal consistent videos, surpassing the performance of ImageDream. For more details on our dataset, please refer to supplementary material.
Training Objectives. For each animated 3D shape from our dataset, we choose a monocular video with a random viewpoint and four videos with orthogonal viewpoints and select frames from each video at a stride of 4 to create our multi-view video dataset . Here, and represent the video clips from and . is the text prompt captioned by Cap3D [33]. Then, we use following the diffusion loss to train 4DM,
here, denote the image embedding noised from a random noise with a diffuse timestep . is 4DM model parametrized by .
@@
The entire set of trainable parameters is denoted as .
4D-aware SDS. Once 4DM is trained, we employ 4D-aware SDS loss to guide the optimization of our 4D representation. To be concrete, we utilize Eqn. 8 to render four frames video from four orthogonal viewpoints and timestamps. Our 4D-aware SDS injects Gaussian noise into and passes to our multi-view video diffusion to provide gradients to update ,
where denotes the pseudo ground truth denoised from 4DM with the input video as condition. Here, we replace original -based SDS loss with -reconstruction loss as in [44].
Anchor Loss. Accurately estimating the elevation and azimuth of input monocular video within the canonical coordination poses a challenge, thereby hindering direct utilization of the input video for supervising the 4D optimization process. Therefore, we utilize 4DM to produce four orthogonal videos conditioned on the input video and choose one anchor video from the four videos, which holds a viewpoint closest to the viewpoint of the input video. Then, we devise an anchor loss based on the anchor video to enhance the appearance details and facilitate the learning of dynamic NeRF. Since it is challenging for 4DM to ensure pixel-to-pixel alignment of the anchor video, we follow [10] to use image-level perceptual loss, i.e., LPIPS [64] and SSIM, for dynamic NeRF optimization,
where and represent the rendered images and anchor images, is the loss weight. Consequently, our total loss function for 4D generation is,
@@ 4.1 Comparisons on 4D Generation
Qualitative Evaluation. To validate 4Diffusion for 4D generation, we compare it to Consistent4D [19], DreamGaussian4D [42], and 4D-fy [4] on monocular video-to-4D task. We illustrate the results in Fig. 4. 4D-fy [4] achieves state-of-the-art results on text-to-4D task and can be simply extended to video-prompt 4D generation by replacing 3D-aware diffusion model. Here, we utilize ImageDream as the 3D-aware diffusion model in 4D-fy. 4D-fy produces 4D contents with inconsistent temporal appearance, sometimes diverging significantly from the input video, as depicted in the first two columns of Fig. 4. This is primarily because integrating gradients from multiple diffusion models is difficult and they face challenges in multi-view spatial-temporal modeling. Consistent4D is the first work for 4D generation from monocular video. They employ an interpolation loss between two frames to enhance spatial-temporal consistency. However, they lack temporal consistency across frames, leading to poor appearance quality and flickers. DreamGaussian4D generates 4D contents with a blurred appearance and inaccurate geometry because GS struggles to model thin structures and large motions under unconstrained situations. In contrast, 4Diffusion generates high-quality 4D content with 4DM, which captures multi-view spatial-temporal correlations in a unified manner. Overall, our method achieves superior results, demonstrating its effectiveness.
Quantitative Evaluation. We employ datasets from ImageDream to evaluate 4D generation results. To evaluate image quality, we leverage CLIP-I [40] to measure the similarity between frames in synthetic videos and the input video. Moreover, we compute CLIP-C between frames in each synthetic video to evaluate temporal consistency. Moreover, we calculate FVD between the input video and synthesized videos to evaluate the video quality. Tab. 1 presents the results, clearly demonstrating that 4Diffusion outperforms other methods on all metrics.
| Image quality | Tem. Con. | Video quality | |
| CLIP-I | CLIP-C | FVD | |
| Consistent4D[19] | 0.8713 | 0.9641 | 1246.9 |
| DreamGaussian4D[42] | 0.8587 | 0.9625 | 1451.2 |
| 4D-fy[4] | 0.8287 | 0.9379 | 2283.6 |
| Ours(w/o ) | 0.7945 | 0.9423 | 2103.2 |
| Ours(w/o ) | 0.8518 | 0.9557 | 1647.0 |
| Ours | 0.8803 | 0.9654 | 1196.8 |
@@ 4.2 Multi-view Video Generation
Qualitative Evaluation. In this section, we evaluate the multi-view video generation quality produced by 4DM. We employ ImageDream to synthesize a set of multi-view images as pseudo multi-view video by taking each frame of the input monocular video as an image prompt. Fig. 5 illustrates results with the first two columns corresponding to the input monocular video. Although ImageDream excels at synthesizing spatially consistent images, it struggles to model temporal correlations, leading to inconsistent temporal appearances, such as the icon on the back of Spiderman. Comparatively, 4DM effectively captures reasonable temporal information using the motion module, even when trained on a small curated dataset. Moreover, our model preserves the generalizability of ImageDream, allowing us to generate high-fidelity multi-view videos, even beyond the distribution of our training dataset. As the last two rows of Fig. 5 show, 4DM produces spatially consistent videos by sharing information across spatial and temporal dimensions to constraint the generation process while ImageDream occasionally fails to generate videos coherent to the viewpoint.
Quantitative Evaluation. We employ Consistent4D test dataset, which contains a monocular video as input and four orthogonal videos as ground truth, to evaluate our proposed model. We leverage the same metrics in Sec. 4.1 and calculate metrics between synthesized videos and ground truth. Additionally, we calculate LPIPS [64] between ground truth and synthesized videos from the same viewpoint to evaluate the spatial consistency. Tab. 2 shows comparative results. 4DM achieves a higher LPIPS, CLIP-C, and FVD, along with comparable CLIP-I compared to ImageDream, reflecting that 4DM efficiently captures spatial-temporal correlations.
| Image quality | Spa. Con. | Tem. Con. | Video Quality | |
| CLIP-I | LPIPS | CLIP-C | FVD | |
| ImageDream[52] | 0.8776 | 0.1704 | 0.9412 | 591.25 |
| Ours(w/ whole) | 0.8365 | 0.2253 | 0.9637 | 889.44 |
| Ours(4DM) | 0.8630 | 0.1644 | 0.9640 | 523.90 |
@@ 4.3 Ablation study and analysis
Effectiveness of the selected multi-view video dataset. To evaluate the importance and effectiveness of the selected high-quality multi-view videos, we use the entire animated 3D shapes from Objaverse and render multi-view videos to fine-tune 4DM (Ours w/ whole). The results are shown in Tab. 2. Given the presence of numerous flawed data within the entire dataset, it compromises the image quality of ImageDream and encounters challenges in precisely capturing spatial-temporal correlations, demonstrating the importance of high-quality datasets for fine-tuning 4DM.
4D-aware SDS Loss. To evaluate the effect of our 4D-aware SDS loss, we substitute the 4DM with ImageDream and use 3D-aware SDS loss based on ImageDream to optimize dynamic NeRF representation. As Fig. 6 depicted, inconsistent temporal textures, such as the leg of squirrel, emerge due to the lack of temporal modeling of ImageDream, underscoring the significance of capturing spatial-temporal corelations in 4DM. The quantitative results presented in Tab. 1 indicate the significance of 4DM.
Anchor Loss. We also assess the impact of the proposed anchor loss. As illustrated in Fig. 6, capturing detailed appearance features, such as the eyes of the squirrel, proves challenging without the anchor loss. Conversely, the anchor images furnish visual clues to facilitate the learning of 4D representation, resulting in high-quality 4D content. The quantitative results Tab. 1 demonstrate the crucial role of the anchor loss.

### v2

@@ 1 Introduction
In this paper, we present a novel 4D generation pipeline, namely 4Diffusion, to create high-quality spatial-temporally consistent 4D content from a monocular video. Specifically, we propose a unified diffusion model, 4DM, to capture multi-view spatial-temporal correlations for multi-view video generation. To achieve this, we construct 4DM based on the powerful pre-trained 3D-aware diffusion model [52], which already ensures high-quality multi-view spatial consistency. We then seamlessly integrate a motion module into the 3D-aware diffusion model to extend the temporal modeling capability. Contrary to previous attempts [16, 14] that typically demand extensive large-scale video datasets for tuning the motion module, 4DM achieves reasonable temporal consistency and captures multi-view spatial-temporal correlations after training on only hundreds of multi-view videos. Importantly, we keep the parameters of the 3D-aware diffusion model unchanged to preserve the generalization ability and spatial consistency. 4DM provides multi-view spatial-temporal guidance for 4D generation. Therefore, we propose 4D-aware SDS loss to distill prior knowledge from 4DM to optimize 4D content parameterized by dynamic NeRF. This approach eliminates discrepancies arising from multiple diffusion models and stabilizes the optimizing process. Moreover, we use 4DM to generate anchor videos conditioned on the input monocular video and devise an anchor loss to enhance the appearance details, facilitating the learning of dynamic NeRF. Finally, we generate 4D content with high-quality spatial-temporal consistency and vibrant motion coherence with the input video as shown in Fig. 1. Qualitative and quantitative experiments demonstrate that our method achieves state-of-the-art performance on multi-view video generation and 4D generation from monocular videos.
@@ 2 Related Work
Similar to us, [20, 42, 63, 62, 57] generate 4D content from a monocular video. Consistent4D [20] introduces an interpolation-driven loss between two adjacent frames to enhance spatial-temporal consistency. However, Consistent4D lacks temporal modeling cross frames. DreamGaussian4D [42], 4DGen [62], and SC4D [57] combine 4D Gaussian Splatting [56, 19] into 4D generation pipeline. Although they notably reduce optimization time, they may result in blurred appearance and inaccurate geometry due to the explicit characteristics of Gaussians. STAG4D [63] proposes a training-free strategy to generate sparse anchor multi-view videos for 4D generation. In contrast, we propose a multi-view video diffusion model to provide multi-view spatial-temporal consistency guidance for 4D generation.
@@ 3 Method
Given a monocular video with frames and an optional textual caption, our goal is to generate a high-quality spatial-temporally consistent 4D content, capable of rendering from any novel viewpoint across the temporal dimension. In Sec. 3.1, we talk about 3D-aware diffusion models, employed as the initialization of our unified diffusion model. In Sec. 3.2, we propose a unified diffusion model 4DM to capture multi-view spatial-temporal correlations for multi-view video generation. Subsequently, we elaborate on distilling prior knowledge from 4DM to optimize 4D content parameterized by dynamic NeRF and devise an anchor loss to enhance the appearance details, as detailed in Sec. 3.3. Fig. 3 shows the overall pipeline of our method.
3.1 Preliminary: 3D-aware Diffusion Models
@@
Although we can directly use the original ImageDream to generate a set of individual multi-view images to form multi-view videos, the result lacks temporal consistency as ImageDream has no layer for temporal modeling, as shown in Fig. 8. We thus add a zero-initialized motion module at the end of each block of the UViT network of ImageDream. Specifically, each motion module begins with group normalization and a linear projection, followed by two self-attention blocks and one feed-forward block. A final linear projection is then applied, after which the residual hidden feature is added back at the end of each motion module as detailed in Fig. 5. Then, each attention block of 4DM includes a spatial module and a motion module . The spatial module comprises a 3D self-attention module and a cross-attention module. We first concatenate the monocular video latent and four multi-view video latents encoded by VAE [22] to obtain a batch of latents , where is the number of channels, and are spatial resolutions, is the number of viewpoints, and is the number of frames. Subsequently, we reshape the temporal axis into the batch dimension and independently process multi-view video latents through the 3D self-attention module,
@@
We utilize Objaverse dataset [11] to train 4DM. Although Objaverse provides nearly 44K animated 3D shapes, rendering multi-view videos and training a diffusion model are time- and computation-consuming using the entire dataset. Moreover, it is worth noting that the Objaverse dataset contains a significant amount of flawed data. Consequently, we manually select a curated subset of 926 high-quality animated 3D shapes from Objaverse dataset [11]. We render multi-view videos from those animated 3D shapes to tune our motion module while holding the parameters of the origin ImageDream frozen. Surprisingly, 4DM successfully learns reasonable temporal dynamics and preserves the characteristics of the origin ImageDream model, including generalization ability, spatial consistency, and image understanding ability, even when trained on a small curated dataset. As Fig. 8 illustrates, 4DM generates multi-view spatial-temporal consistent videos, surpassing the performance of ImageDream. For more details on our dataset, please refer to supplementary material.
Training Objectives. For each animated 3D shape from our dataset, we render a monocular video with a random viewpoint and four videos with orthogonal viewpoints and select frames from each video at a stride of 4 to create our multi-view video dataset . Here, and represent the video clips from and . is the text prompt captioned by Cap3D [33]. Then, we use following the diffusion loss to train 4DM,
here, represent the noisy video latents derived from . These latents are initially encoded by VAE and subsequently noised by random noise at a diffusion timestep . For more details about the noising process, please refer to [43]. is 4DM model parametrized by . During the training of 4DM, we ensure that our training data does not overlap with the test data used in our experiments.
@@
The entire set of trainable parameters is denoted as . We can optimize our dynamic NeRF by using the multi-view videos generated from 4DM, however, 4DM can only produce four orthogonal viewpoints at one time. Training with such sparse views often results in overfitting to the training viewpoints, as presented in Fig. 6. To mitigate this, we leverage 4D-aware SDS to optimize the dynamic NeRF, enabling effective rendering from novel viewpoints across the temporal dimension, which is crucial for 4D generation.
4D-aware SDS. Once 4DM is trained, we employ 4D-aware SDS loss to guide the optimization of our 4D representation. To be concrete, we utilize Eqn. 8 to render four frames video with timestamps from four orthogonal viewpoints . Our 4D-aware SDS injects Gaussian noise into at a diffusion timestep and passes to our multi-view video diffusion to provide gradients to update ,
where denotes the pseudo ground truth denoised from 4DM with the input video and viewpoints as condition. Here, we replace original -based SDS loss with -reconstruction loss as in [44].
Anchor Loss. Accurately estimating the elevation and azimuth of input monocular video within the canonical coordination is challenging, making it difficult to use the input video directly as supervision signals. Therefore, we utilize 4DM to produce four orthogonal videos conditioned on the input video and select the one with the viewpoint closest to that of the input video as the anchor video. This approach ensures that the anchor video maintains the same quality as the input and improves the results. Moreover, 4DM is currently limited to generating multi-view videos with 8 frames. When the input video exceeds 8 frames, we must apply our multi-view video diffusion model multiple times to generate anchor videos. However, this process may lead to temporally inconsistent results due to the stochasticity of the diffusion model, particularly when the viewpoint is far from the input video as shown in Fig. 5. This inconsistency would degrade the 4D generation performance. Finally, we devise an anchor loss based on the anchor video to enhance the appearance details and facilitate the learning of dynamic NeRF. Since it is challenging for 4DM to ensure pixel-to-pixel alignment of the anchor video, we follow [10] to use image-level perceptual loss, i.e., LPIPS [64] and SSIM, for dynamic NeRF optimization,
where and represent the rendered video and anchor video, is the loss weight. Consequently, our total loss function for 4D generation is,
@@ 4.1 Comparisons on 4D Generation
Qualitative Evaluation. To validate 4Diffusion for 4D generation, we compare it to Consistent4D [20], DreamGaussian4D [42], and 4D-fy [4] on monocular video-to-4D task. Here, we use 3 real-world videos and 3 synthetic videos from the Consistent4D dataset, as well as 3 images from the ImageDream. As discussed in Sec.A.1 of the supplementary materials, for text-image pairs from ImageDream, we utilize SVD to generate input videos. We illustrate the results in Fig. 7. 4D-fy [4] achieves state-of-the-art results on text-to-4D task and can be simply extended to video-prompt 4D generation by replacing 3D-aware diffusion model. Here, we utilize ImageDream as the 3D-aware diffusion model in 4D-fy. 4D-fy produces 4D contents with inconsistent temporal appearance, sometimes diverging significantly from the input video, as depicted in the first two columns of Fig. 7. This is primarily because integrating gradients from multiple diffusion models is difficult and they face challenges in multi-view spatial-temporal modeling. Consistent4D is the first work for 4D generation from monocular video. They employ an interpolation loss between two frames to enhance spatial-temporal consistency. However, they lack temporal consistency across frames, leading to poor appearance quality and flickers. DreamGaussian4D generates 4D contents with a blurred appearance and inaccurate geometry because GS struggles to model thin structures and large motions under unconstrained situations. In contrast, 4Diffusion generates high-quality 4D content with 4DM, which captures multi-view spatial-temporal correlations in a unified manner. Overall, our method achieves superior results, demonstrating its effectiveness. For more visualization results, please refer to our supplementary materials.
Quantitative Evaluation. We select 5 test cases from Objaverse, each consisting of a monocular input video and four orthogonal ground truth videos, which are not included in the training data, to evaluate our model. To evaluate image quality, we leverage CLIP-I [40] to measure the similarity. We also calculate FVD to evaluate the video quality. We compute LPIPS [64] and PSNR metrics to evaluate the spatial consistency. Here, we use ground truth videos for novel viewpoints to compute the above metrics. Moreover, we compute CLIP-C between frames in each synthetic video to evaluate temporal consistency. Tab. 1 presents the results, clearly demonstrating that 4Diffusion outperforms other methods on all metrics.
Image quality | Tem. Con. | Video quality | Spa. Con. | ||
CLIP-I | CLIP-C | FVD | LPIPS | PSNR | |
4D-fy[4] | 0.8658 | 0.9487 | 1042.3 | 0.2254 | 14.24 |
Consistent4D[20] | 0.9216 | 0.9723 | 706.07 | 0.1593 | 16.70 |
DreamGaussian4D[42] | 0.8898 | 0.9710 | 760.18 | 0.1793 | 15.97 |
Ours(w/o ) | 0.8195 | 0.9503 | 1546.4 | 0.2356 | 13.92 |
Ours(w/o ) | 0.8823 | 0.9720 | 853.57 | 0.1589 | 17.20 |
Ours | 0.9310 | 0.9798 | 417.63 | 0.1199 | 19.07 |
@@ 4.2 Multi-view Video Generation
Qualitative Evaluation. In this section, we evaluate the multi-view video generation quality produced by 4DM using the same input videos as described in qualitative evaluation in Sec 4.1. We employ ImageDream to synthesize a set of multi-view images as pseudo multi-view video by taking each frame of the input monocular video as an image prompt. Fig. 8 illustrates results with the first two columns corresponding to the input monocular video. Although ImageDream excels at synthesizing spatially consistent images, it struggles to model temporal correlations, leading to inconsistent temporal appearances, such as the icon on the back of Spiderman. Comparatively, 4DM effectively captures reasonable temporal information using the motion module, even when trained on a small curated dataset. Moreover, our model preserves the generalization ability of ImageDream, allowing us to generate high-fidelity multi-view videos, even beyond the distribution of our training dataset. As the last two rows of Fig. 8 show, 4DM produces spatially consistent videos by sharing information across spatial and temporal dimensions to constraint the generation process while ImageDream occasionally fails to generate videos coherent to the viewpoint.
Quantitative Evaluation. We use the same test cases described in quantitative evaluation in Sec 4.1, alongside the test data provided by Consistent4D, to evaluate 4DM. To account for the stochasticity of the diffusion model, we conduct five runs for each test case and report the average metrics. Tab. 2 shows comparative results. Despite the comparable performance in CLIP-I, 4DM excels in generating spatial-temporally consistent multi-view videos, a primary focus of our research. This is evidenced by the superior performance on metrics such as CLIP-C, FVD, LPIPS, and PSNR, which better capture the spatial and temporal fidelity of video content. These metrics demonstrate that our method effectively balances image quality with temporal consistency, making it a robust solution for multi-view video generation.
Image quality | Tem. Con. | Video Quality | Spa. Con. | ||
CLIP-I | CLIP-C | FVD | LPIPS | PSNR | |
ImageDream[52] | 0.9165 | 0.9320 | 465.94 | 0.1536 | 16.57 |
Ours(w/ whole) | 0.8872 | 0.9478 | 583.79 | 0.1763 | 15.28 |
Ours(4DM) | 0.9260 | 0.9601 | 427.34 | 0.1346 | 17.88 |
@@ 4.3 Ablation study and analysis
Effectiveness of the Curated Multi-view Video Dataset. To evaluate the importance and effectiveness of the selected high-quality multi-view videos, we use the entire animated 3D shapes from Objaverse and render multi-view videos to fine-tune 4DM (Ours w/ whole). The results are shown in Tab. 2. Given the presence of numerous flawed data within the entire dataset, it compromises the image quality of ImageDream and encounters challenges in precisely capturing spatial-temporal correlations, demonstrating the importance of high-quality datasets for fine-tuning 4DM.
4D-aware SDS Loss. To evaluate the effect of our 4D-aware SDS loss, we substitute the 4DM with ImageDream and use 3D-aware SDS loss based on ImageDream to optimize dynamic NeRF representation. As Fig. 9 depicted, inconsistent temporal textures, such as the leg of the squirrel, emerge due to the lack of temporal modeling of ImageDream, underscoring the significance of capturing spatial-temporal correlations in 4DM. The quantitative results presented in Tab. 1 indicate the significance of our 4D-aware SDS loss.
Anchor Loss. We also assess the impact of the proposed anchor loss. As illustrated in Fig. 9, capturing detailed appearance features, such as the eyes of the squirrel, proves challenging without the anchor loss. Conversely, the anchor images furnish visual clues to facilitate the learning of 4D representation, resulting in high-quality 4D content. The quantitative results Tab. 1 demonstrate the crucial role of our anchor loss.